Is the squence geometric? If so identify the common ratio 6,12, 24, 48,....
Found 2 solutions by amfagge92, Alan3354:
Answer by amfagge92(93)
Answer by Alan3354(69443)

2
------
Is that not obvious?
12/6 = 2
24/12 = 2
48/24 = 2